3- Consider the third example solved in the lecture videos. Two additional 12-ft through lanes are added to Vine Street (the street in the intersection of the first example solved in the lecture videos), one lane in each direction. If the peak-hour traffic volumes are unchanged but the Vine Street left-turn saturation flow rates increase by 120 veh/h because of the added through lanes, what would the revised effective green time, yellow time, and all-red time be for each phase? Assume minimum cycle length and a critical intersection v/c of 0.95 is desired.

4- Consider Problem 3. Calculate the new minimum pedestrian green time, assuming the effective crosswalk width is 6 ft and the maximum number of crossing pedestrians in any phase is 20.
Answer #1

ANSWER : © 390 NBTR = 3.18 NBTR = 0.108. we know that formula Yc =EBL + WBTR + NBL 50.171+0.338 +0.157 Yc = = 0.666 we have L=12 give& 50-666.X50 50-12 33.3 38 XE 0.876 (min o .EBL. بالا go8sec 8, (min WBTR 2 xa -0,338 X 50 0.876 19.3dec or (min NBL IC 0.157we assumed that tr=10, 2=20, a = 10 ft/s2 Whale boft we have, Y vine ert Vuire 009 quine 3.567 Ivine = 40 sec 4 I AR vine = W4 Imaple = txt maple 2.9. = 3.993 20 Imaple = 4.0 sec we know that Almaple Wrine +1 maple 1.364. Armaple = 1.5sec Here given8 70 550 SBL=0.127 SBTR = 370 /2018 SBTR = 0.103 go NBL = = 0.157 ㄷ 90 475 + 100 575 NBL = 0.157